Grade 12th passIIT JEE Entrance Exam

If it 12 out of 30 what is it out of 80?

TEJKIRAT SINGH
11 Years agoGrade 12th pass

2 Answers

Harsh Patodia
ApprovedApproved Tutor Answer11 Years ago
12/30 = x/80
x=12*80/30
x=32

32 out of 80
Vikas TU
ApprovedApproved Tutor Answer11 Years ago
U can do this type of prblm... (unitary method)
like this:
 
12 == 30
y == 50
now cross multiply...
y == 12 x 50 /30 == 60/3 = 20